### v1

@@ Abstract
We describe the heuristics used by the Shadoks team in the CG:SHOP 2024 Challenge. Each instance consists of a convex polygon called container and a multiset of items, where each item is a simple polygon and has an associated value. The goal is to pack some of the items inside the container using translations, in order to maximize the sum of their values. Our strategy consists of obtaining good initial solutions and improving them with local search. To obtain the initial solutions we used integer programming and a carefully designed greedy approach.
@@ 1 Introduction
CG:SHOP Challenge is an annual geometric optimization challenge. The sixth edition in 2024 considers a 2-dimensional knapsack packing problem. The team Shadoks won first place with the best solution (among the $14$ participating teams) to $75$ instances out of $180$. In this paper, we describe the heuristics we used. We start by briefly describing the problem.
An input instance consists of a convex polygon called container and a multiset of items. Each item is a simple polygon with an associated integer value. The goal is to pack some of the instance items inside the container using integer translations in order to maximize the sum of their values. In total, $180$ instances have been given, ranging from $28$ to $50{,}000$ items. The instances are of several different types according to the shape and values of the items. Some instances have mostly convex items, while other instances have many non-convex items such as polyominoes. In terms of item values, some items have only unit value, some have values proportionals to the area, and other have random values, for example. Some solutions are presented in Figure 1 and more details about the challenge are available in the organizers’ survey paper [3].
Our general strategy consists of finding a good initial solutions (using integer programming or a greedy heuristic) and subsequently optimizing them with local search. Our strategy shares many common elements with the second place [5], but they did not use integer programming to obtain initial solutions and their optimization phase is more sophisticated than ours. The third place [4] uses an hierarchical grid approach. The fourth place [1] used a completely different integer programming model and a genetic algorithm.
We describe the algorithms in Section 2 and experimentally analyze their performance using different parameters in Section 3. Our solvers were coded in Python and C++ and executed on several desktop laptop computers, as well as the LIMOS and LIS clusters.
@@ 2.1 Integer Programming Approach
A simple idea to solve the challenge problem is to produce a set $V$ of random translations of each item inside the container and then reduce the problem to a kind of maximum weight independent set problem in a graph $G=(V,E)$. Each translated item is a vertex and there are two types of edges: (1) an edge between two translations that overlap and (2) translations of the same item $i$ form a clique $C_{i}$. If all item have quantity one, then this is a traditional maximum weight independent set problem. However, if items have non-unit quantities, then each clique $C_{i}$ is associated with the quantity $q_{i}$ of item $i$ and at most $q_{i}$ vertices of the clique are allowed in the solution.
@@
To obtain better solutions, we start from a solutions $S$ obtained with the aforementioned method and build a new graph $G=(V,E)$ as follows. Let $\sigma>0$ be a parameter and $N$ be a set of the zero vector and random vectors where each random vector has $x$ and $y$ coordinates as Gaussian random variables of average 00 and standard deviation $\sigma$. We create a translation in $V$ for each item that is placed in $S$ and for each translation vector in $N$ if the translation is inside the container. We also create vertices in $V$ using uniform random translations for all items. Edges and cliques are created as before, and the new combinatorial problem is solved with CPLEX. We repeat this procedure multiple times using the previous solution $S$ and reducing the value of $\sigma$ at each step.
@@ 2.2 Greedy Heuristic
The greedy heuristic starts by creating an initial list $L$ of $n$ grid points inside the container (typically $n=1000$). The list $L$ is shuffled and we compute its centroid $c$ rounded to integer coordinates. The point $c$ is inserted in the beginning of $L$. The input items are placed into a list $I$ ordered by decreasing utility, where the utility function is described next. Different utility functions may be used (more details in Section 3). The goal is that items of small area and high value have high utility while large items with small value have low utility.
We then build the packing by considering the items of the list $I$ one by one in order. At each step, we have a current packing and a new item $i$ to pack. We first try to pack $i$ at the first grid position $g\in L$ and at a number of random positions around $g$. If one of these positions is available, i.e the item is fully contained in the container and does not cross any currently packed item, then we pack item $i$ at the corresponding position and proceed to the next item in $I$. Otherwise, we go to the next position. If all the positions have been tried without success, then we go to the next item in $I$.
Immediately after an item $i\in I$ is packed, its position is modified as follows. We choose a direction $u$ in which $i$ is pushed until it hits another item or the container boundary. Pushing an item means that we translate it in the direction $u$ and in directions $v$ where the dot product of $u$ and $v$ is positive until no improvement in any direction $u$ or $v$ can be achieved.
The greedy algorithms ends when all items in $I$ have either been packed or tested at all the positions of the list $L$.
2.3 Slates Preprocessing
The goal of the preprocessing step is to precompute some sets of items that assemble together efficiently. We call slate a set of items with a fixed relative position. A slate plays the same role as an item, but it is a polygonal set instead of a simple polygon. As computing slates takes a lot of time, we control it by building graphs of items. The goal of these graphs is to tell the pairs of items that we want to try to assemble together. A first graph contains for instance all the thin items and connects the items with similar diameter directions. Another graph contains the non convex items and connects them with items that can fill their concavities. We also compute graphs containing hundreds of random items and connecting items having a close pair of edges $uv$ and $u^{\prime}v^{\prime}$ where the Euclidean distance $d(v-u,u^{\prime}-v^{\prime})$ is small compared to $d(u,v)$. For each graph, we try to assemble adjacent items by searching for a relative position minimizing the area of the convex hull of their union. Each slate composed by two items has a utility and we only keep the slates of high utility. We take the slates composed by two items and try to assemble a third adjacent item of the graph. We get a new generation of slates with $3$ items and continue the growing process until a fixed level. Again at each generation, we keep only the slates of high utility.
@@ 2.3 Slates Preprocessing
Figure 2: Slates for the instances jigsaw_cf2_xf42cb20_670, random_cf3_x21f5def_200, satris1685, and atris1660. The items of a slate are drawn with the same color.
Slates work particularly well for the atris instances where the items resemble polyominoes.
@@ 2.4 Local Search
To improve the quality of the solutions obtained with the previous approaches, we use a local search optimization scheme by repeating the following routine until a time limit is reached. We randomly choose between executing two routines, either a fill routine or a push routine. The fill routine is similar to the greedy heuristic. We try to pack all the unpacked items at several grid and random positions.
The push routine first randomly chooses either a point $v$ in the interior of the container or a vertex of the container. The packed items are sorted by the distance of their centroids to the point $v$, from the farthest to the closest. Then, each packed item $i$ with centroid $c_{i}$ is pushed in the direction $c_{i}-v$. The goal is to free the space around $v$. Once all packed items have been pushed, we try to pack the unpacked items around $v$ by using the grid points $g$ in $L$ close to $v$ and some random integer points around each point $g$. It may be the case that an item can be packed if we remove some other items of the packing that intersect it. In this case, we compute the benefit of removing the conflicting items in order to pack the new item, replacing the item if the benefit is non-negative. To accelerate the procedure for large instances, we add a parameter which allows us to push only the items that are close to $v$.
@@ 3 Parameter Analysis
In order to compare the different parameters, we consider $18$ instances of various sizes, as shown in Table 1. Throughout, we refer to the ratio between the value of the solution in question and the value of the best solution in the challenge as the value ratio (notice that the competition score is the square of the value ratio). The best solutions to the first $15$ table instances are shown in Figure 3.
@@
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 4 shows the evolution of the value ratio by using the slate preprocessing step, the greedy algorithm followed by local search until it becomes stable with standard parameters for three instances with $100$, $200$ and $1000$ items, respectively. It shows that sometimes the slates provide a benefit, while sometimes it has a negative impact on the results.
random_cf4_50e0d4d9_100
random_cf3_x21f5def_200
random_rcf3_x7651267_1000
Figure 4: The value ratio over time for the instances random_cf4_50e0d4d9_100, random_cf3_x21f5def_200, and random_rcf3_x7651267_1000 for $12$ independent executions of the greedy algorithm and a subsequent local search. Red curves are computed without slate preprocessing, while the light blue and dark blue curves respectively include a preprocessing of slates with at most $5$ and $10$ items.
We now analyze the parameters of our algorithms. The greedy algorithm depends on several parameters:
  • The number of grid positions and the number of random integer positions to consider around each of these points. We typically used numbers from $500$ to $3000$ grid positions plus from $3$ to $10$ random positions around each grid point. Figure 5 compares different choices of the parameter and shows that there is no major improvement to wait for by using more than $1000$ grid positions.
  • The utility function used to sort $L$ may be the value of the item, the value of the item divided by its area, the value of the item to a power $1.5$ divided by its area with or without complementary weights. For the Table 1 instances, we show in Figure 6 the value ratio first after greedy and second after optimization for different utility functions.
  • The direction $u$ to push an item may be chosen randomly (strategy 1), or as a normal to the item diameter (with a random choice to determine which side, strategy 2). However, we mainly used the following direction choice. If the item is thin (ratio between its diameter and its width normal to the diameter above $3$), we push it in the normal direction to its diameter but to the left while if it is fat (ratio below $3$) we push it to the right (strategy 3). Finally, we may push the items in a direction normal to their longest edge if they are fat (strategy 4). The value of the average for 10 runs on each Table 1 instance is shown in Figure 7.

### v2

@@ Abstract
The 2024 edition of the CG:SHOP Challenge focused on the knapsack polygonal packing problem. Each instance consists of a convex polygon known as the container and a multiset of items, where each item is a simple polygon with an associated integer value. A feasible packing solution places a selection of the items inside the container without overlapping and using only translations. The goal is to achieve a packing that maximizes the total value of the items in the solution.
Our approach to win first place is divided into two main steps. First, we generate promising initial solutions using two strategies: one based on integer linear programming and the other on employing a combination of geometric greedy heuristics. In the second step, we enhance these solutions through local search techniques, which involve repositioning items and exploring potential replacements to improve the total value of the packing.
@@ 1 Introduction
The CG:SHOP Challenge is an annual competition in geometric optimization. In its sixth edition in 2024, the challenge focuses on a two-dimensional knapsack packing problem. The team Shadoks won first place with the best solution (among the 14 participating teams) to $75$ instances out of $180$ instances.
In this paper, we outline the heuristics we employed, beginning with a brief overview of the problem. Each input instance comprises a convex polygon referred to as the container and a multiset of items, where each item is a simple polygon associated with an integer value. The objective is to pack a selection of the items inside the container using integer translations, maximizing the total sum of their values.
A total of 180 instances were provided, ranging from $28$ to $50{,}000$ items. The instances are categorized into four classes: Atris, Satris, Jigsaw, and Random, based on the shapes of the items. The Jigsaw instances contain convex polygons with edges in random directions as items and convex containers. The Random instances contain arbitrary polygons as items and convex containers. The Atris instances contain polyominoes as items and rectangular containers. The Satris instances contain polygons obtained by shearing polyominoes as items and rectangular containers. The item values also vary. Some instances have all items with unit value, other instances have values that are small random integers, and other instances have values that are roughly proportional to the area. Solutions to an instance of each class are illustrated in Figure 1, and additional details about the challenge can be found in the organizers’ survey paper [3].
Our general strategy consists of finding good initial solutions (using integer programming or a greedy heuristic) and subsequently optimizing them with local search. Our strategy shares a geometric greedy approach with the second-place team [5], but they did not use integer programming to obtain initial solutions, and their optimization phase is also different from ours. The third-place team [4] uses a hierarchical grid approach. The fourth-place team [1] employed a completely different integer programming model and a genetic algorithm.
We describe the algorithms in Section 2 and experimentally analyze their performance using different parameters in Section 3. Our solvers were coded in Python and C++ and executed on several desktop and laptop computers, as well as the LIMOS and LIS clusters.
@@ 2.1 Integer Programming Approach
A simple idea to solve the challenge problem is to produce a set $V$ of random translations of each item inside the container and then reduce the problem to a kind of maximum weight independent set problem in a graph $G=(V,E)$. Each translated item is a vertex with vertex weights determined by the values of the items. There are two types of edges: (1) an edge between two translations that overlap and (2) translations of the same item $i$ form a clique $C_{i}$. If all item have quantity one, then this is a traditional maximum weight independent set problem. However, if items have non-unit quantities, then each clique $C_{i}$ is associated with the quantity $q_{i}$ of item $i$ and at most $q_{i}$ vertices of the clique are allowed in the solution.
@@
To obtain better solutions, we start from a solutions $S$ obtained with the aforementioned method and build a graph $G=(V,E)$ using the same edge rules as before, but for a different set of vertices $V$ constructed as follows. Let $\sigma>0$ be a parameter and $N$ be a set of the zero vector and random vectors where each random vector has $x$ and $y$ coordinates as Gaussian random variables of average 00 and standard deviation $\sigma$. For each item that is placed in $S$ and for each translation vector in $N$, we create a vertex in $V$. Translations that are not completely inside the container are removed. We also create vertices in $V$ using uniform random translations for all items. Edges and cliques are created as before, and the new combinatorial problem is solved with CPLEX. We repeat this procedure multiple times using the previous solution $S$ and reducing the value of $\sigma$ at each step.
@@ 2.2 Greedy Heuristic
The greedy heuristic begins by generating an initial list $L$ of $n$ grid points within the container, where typically $50001000<n<$ 5000. The list $L$ is then shuffled, and its centroid $c$ is computed and rounded to integer coordinates. The point $c$ is subsequently inserted at the beginning of $L$.
Input items are arranged in a list $I$, sorted by decreasing utility. The utility function, detailed in Section 3, is designed to prioritize items with small area and high value. Some example of utility functions we used are the item’s value, the item’s value divided by its area, or the item’s value raised to the power of $1.5$ divided by its area. By starting with high-utility items, the algorithm aims to maximize value while minimizing occupied area. Although this strategy seems intuitive, alternative approaches were explored, such as placing small items at the end of the list to fill remaining spaces as the container becomes full or during local search. However, determining the threshold size for this approach proved challenging, and the idea was ultimately abandoned due to the challenge time constraints.
Regardless of the sorting criteria used, the packing is then constructed by considering items from the list $I$ one by one in order.
At each step, we have a current packing and a new item $i$ to pack. We first try to pack $i$ at the first position $g\in L$, namely in the center of the container. If it overlaps a previously packed item, we try some random positions around $g$. If we do not find a valid position for item around $g$, we move to the next position $g$ in the list $L$ and try some random positions around it. Then the routine that we repeat at each position $g$ of $L$ is the following: first try to pack item $i$ at the grid position $g\in L$. If there are overlaps with the previously packed items or the boundary of the container, try some random positions around $g$. If no valid positions is found for item $i$ around $g$, skip to the next position.
If all positions in $L$ have been considered without success for item $i$, we skip to the next item in $I$.
When an available position is found for an item $i$, the item placed at this position is fully contained in the container and does not overlap any previously packed item. However, positioning item $i$ at this first available position is not necessarily optimal. We attempt to move item $i$ so that it does not remain in the middle of the empty space in the container.
We select a direction $u$ along which the item $i$ is pushed until it can no longer advance due to obstructions from either other packed items or the boundaries of the container. Pushing an item entails translating it in directions $v$ such that the dot product $v.u$ with $u$ is strictly positive. Since the dot product of the item’s position with $u$ always increases during the process, the possibility of loops is avoided. The push routine terminates when no further movement is possible in any direction $v$.
There are several strategies for selecting the push direction $u$ (see Section 3 for details). The most efficient approach is to push the items in a direction normal to their diameter. Alternatively, a separation criterion can be introduced, where thinner items are pushed to the left and rounder items to the right. The greedy algorithm ends when all items in $I$ have either been packed or tested at all positions in the list $L$. Four distinct strategies are illustrated in Figure 2.
(a) Value ratio: $0.82$                  (b) Value ratio: $0.85$
@@
  (c) Value ratio: $0.89$                   (d) Value ratio: $0.92$
Figure 2: Four solutions for the instance jigsaw_rcf4_6de1b3b7_1363 obtained using the greedy algorithm with different strategies to select the direction $u$ to push the items and their value ratios compared to the best solution found during the challenge. (a) The direction $u$ is randomly chosen and fixed for all items. (b) $u$ is set to be normal to each item’s diameter with negative $y$ coordinate. (c) $u$ is again normal to the item’s diameter, but with a random choice of left or right. (d) $u$ is normal to the diameter and directed left for skinny items and right for fat items.
2.3 Clusters Preprocessing
Our greedy heuristic does not mimic the human approach to solve small packing instances. When faced with a packing puzzle, a human would likely first group items into compact clusters before attempting to pack them into the container. In other words, a human would preprocess the items to identify useful clusters.
The objective of the preprocessing step is to combine small sets of items into larger groups. The term cluster refers to a set of items with fixed relative positions. A cluster functions like an individual item but is a polygonal set rather than a simple polygon.
Problem statement.
Computing clusters is an ill-posed problem until we establish an objective function to define a ’good’ cluster. Utility functions from greedy algorithms could serve as candidates, but clusters introduce empty space between items, reducing their utility compared to individual items. For instance, if the utility function is the ratio between the sum of values and the convex hull area, the cluster’s utility is often smaller than the utility of its constituent items. Consequently, clusters may seem less competitive and be deprioritized in packing orders. Thus, the primary aim is to design a cluster utility function to effectively evaluate cluster efficiency.
Given a known efficient packing of the container, the ratio of the total item value to the container area offers insight into the minimum value density required for improvement. This ratio serves as a threshold to reject low-density clusters (and items). To encourage large cluster usage, we make the cluster utility superlinear in value:
cluster utility($\text{cluster utility}(cluster)=\frac{gauss\cdot value^{\alpha}\cdot penalty}{area}$
where
  • $gauss$ is a random Gaussian variable with mean $1$ to diversify clusters upon repeated computation,
  • $value$ is the sum of item values,
  • $\alpha$ is an exponent between $1$ and $2$,
  • $area$ is either the convex hull area or a weighted sum with items areas,
  • $penalty$ is fixed equal to $1$ except in some cases presenting clusters that we would like to discourage. Let us consider for instance that we have $10$ copies of an axis-parallel rectangular of width $10$ and height $1$. There are two ways to assemble the copies of this item: either in a long bar of size $100\times 1$ or in a square of size $10\times 10$. The long bar does not seem a good choice but without penalty, its cluster utility would be equal to the one of the square. Then we penalize thin clusters from a factor $penalty<1$ computed by using the thickness of the items and the thickness of the cluster.
The preprocessing goal is to compute clusters with high cluster utility.
General strategy.
Some technical details depend on item types (polyominoes, sheared polyominoes, convex, non-convex) but we first outline the common pipeline before delving into specific shape classes.
With thousands of items, evaluating all possible pairs for assembly is infeasible. Thus, the first task is to construct a compatibility graph connecting items that can potentially be assembled together. But instead of computing a unique compatibility graph, we compute several according different criterion. It provides a better control on the clusters generation.
The second step consists in generating clusters from a compatibility graph. We start by generating clusters with $2$ items until a fixed maximum number of items. The items are the generation $1$ and the generation $k+1$ (clusters with $k+1$ items) is built from the generations $1$ and $k$. The clusters of the generation $k+1$ items are computed by assembling clusters of the generation $k$ with items which are compatible to at least one of their own items. This implies that the items within a cluster of generation $k$ form a connected component of size $k$ in the compatibility graphs. If the average degree of the compatibility graphs is larger than a small constant, it becomes impractical to attempt all possibilities. A crucial requirement is to manage the exponential growth of the number of potential clusters. We achieve this goal in all generations by keeping only a fixed number $m$ of clusters per item. The selection is done according to the cluster utility function. For each item, we keep the $m$ clusters of highest cluster utility.
Building the compatibility graphs.
If the number of items is small, then we use a unique compatibility graph which is a clique on all items. If the number of items is large, a quadratic number of compatibilities is too large. Instead we compute several compatibility graphs of smaller size:
  • The first compatibility graph is denoted Rand. We first define a partition of the set of items in subsets of fixed size (e.g., $100$). The compatibility graph $Rand$ is the union of the cliques of this partition subsets.
  • A second compatibility graph is denoted Thin. This graph connects the items whose ratio diameter/width is larger than a constant (e.g., $3$) and whose longest edges have roughly the same direction.
  • A third graph is called Concav. It connects some items with concavities with some items whose shape could allow us to fill the holes.
The three previous graphs are considered for items with edges in arbitrary directions. For specific instances like Atris and Satris, other graphs can be computed.
  • Satris items are sheared polyominoes. We build a compatibility graph called Shear by taking account the directions of the edges of the items. Each item has two dominant directions. We group the items with close dominant directions and take as Shear the union of the cliques of these groups.
  • Atris items are slightly pertubed polyominoes but the perturbations are sufficiently small to allow us to encode each shape by its Freeman code. Then the complementarity of item pairs can easily checked by word combinatoric. Nevertheless, the instances of Atris used only a small number of shapes (bars, crosses, L, shapes in V and waves). It makes possible to achieve a complete classification of the items and thus build a compatibility graph which connects all the items of a given class to all the items of another one. This compatibility graph is denoted Polyo.
Assembly routine.
We employ two distinct procedures to assemble an item into a cluster:
  • Grid-based approach: Center the cluster on a 2D grid containing between $100$ and $1000$ points. Translate the item to each grid position. For each position, check if the item overlaps the cluster. If not, compute the cluster utility of the union of the cluster and the item at that position.
  • Vertex-based approach: Consider all pairs of item and cluster vertices. Translate the item vertex onto the cluster vertex. If there are no overlaps, then compute the new cluster utility.
We may use either or both methods. Ultimately, we retain the best cluster considered throughout the process. The overall process is the following. First, we build the chosen compatibility graphs. For each graph, we generate clusters of $k$ items by using the assembly routine and keeping only the best clusters containing each item.
@@ 2.3 Clusters Preprocessing
Figure 3: clusters for the instances jigsaw_cf2_xf42cb20_670, random_cf3_x21f5def_200, satris1685, and atris1660. The items of a cluster are drawn with the same color.
When the cluster preprocessing step is enabled in our greedy algorithm, we replace the utility function used for sorting items with the cluster utility function, which is then used to sort all available clusters. Note that when a cluster is packed, the number of copies of its items is reduced, potentially making other clusters unavailable.
Unsurprisingly, clusters work particularly well for the atris instances, where items resemble polyominoes. During the CG 2024 competition, the computation time for clusters was not sufficiently controlled, limiting the use of this strategy for large instances. Consequently, we only made moderate use of clusters.
@@ 2.4 Local Search
To improve the quality of solutions produced by the previous approaches, we employ a local search optimization scheme, which iterates the following routine until a specified time limit is reached. At each iteration, one of two routines is randomly selected: a fill routine or a dig routine. The fill routine, akin to a greedy heuristic, attempts to pack all unpacked items into the grid by testing various grid positions and random placements.
The strategy of the dig routine is simply to choose a point $v$ of the container and to push the packed items as far as possible from this point in order to free some space. Then the dig routine first randomly chooses either a point $v$ in the interior of the container or a vertex of the container. The packed items are sorted by the distance of their centroids to the point $v$, from the farthest to the closest. Then, each packed item $i$ with centroid $c_{i}$ is pushed in the direction $c_{i}-v$. The goal is to free the space around $v$. Once all packed items have been pushed, we try to pack the unpacked items around $v$ by using the grid points $g$ in $L$ close to $v$ and some random integer points around each point $g$. It may be the case that an item can be packed if we remove some other items of the packing that intersect it. In this case, we compute the benefit of removing the conflicting items in order to pack the new item, replacing the item if the benefit is non-negative. To accelerate the procedure for large instances, we add a parameter which allows us to push only the items that are close to $v$.
The strategy of the dig routine is straightforward: it selects a point $v$ within the container and attempts to push the packed items away from this point to free up space.
First, the routine randomly chooses the point $v$, either in the interior or at a vertex of the container. Next, the packed items are sorted by the distance of their centroids from $v$, in descending order. Each packed item $i$, with centroid $c_{i}$, is then pushed in the direction of $c_{i}-v$, aiming to clear the space around $v$.
Once the items have been repositioned, the routine attempts to pack the remaining unpacked items around $v$ by using nearby grid points $v$ from the set $L$ and several random integer points around each grid point. In cases where packing a new item requires removing others due to overlap, the routine computes the benefit of removing the conflicting items by comparing their values. If the benefit of replacing them with the new item is non-negative, the replacement is made.
To speed up the procedure for larger instances, a parameter is introduced to restrict the pushing operation to items near $v$, either by defining a radius or by setting a maximum number of items to push.
@@ 3 Parameter Analysis
In order to compare the different parameters, we consider $18$ instances of various sizes, as shown in Table 1. Throughout, we refer to the ratio between the value of the solution in question and the value of the best solution in the challenge as the value ratio (notice that the competition score is the square of the value ratio). The best solutions we found to the first $15$ table instances are shown in Figure 4.
@@
Figure 5 illustrates the evolution of the value ratio when applying the cluster preprocessing step, followed by the greedy algorithm and local search until stabilization, using standard parameters for three instances with $100$, $200$, and $1000$ items, respectively. The results indicate that while clustering can sometimes improve performance, it may also have a negative impact in certain cases.
random_cf4_50e0d4d9_100 Refer to caption random_cf3_x21f5def_200Refer to caption random_rcf3_x7651267_1000
Figure 5: The value ratio over time for the instances random_cf4_50e0d4d9_100, random_cf3_x21f5def_200, and random_rcf3_x7651267_1000 for $12$ independent executions of the greedy algorithm and a subsequent local search. Red curves are computed without cluster preprocessing, while the light blue and dark blue curves respectively include a preprocessing of clusters with at most $5$ and $10$ items.
We now analyze the parameters of the greedy heuristic:
  • The number of grid positions and random integer positions considered around each point varies. Typically, we used between $500$ and $3000$ grid positions, along with $3$ to $10$ random positions per grid point. Figure 6 compares different parameter settings and shows that using more than $1000$ grid positions offers no significant improvement.
  • The utility function used to sort the items $I$ may use various criteria: the item’s value, the item’s value divided by its area, or the item’s value raised to the power of $1.5$ divided by its area, with or without additional weighting factors. For the instances in Table 1, Figure 7 shows the value ratio after the greedy phase and again after optimization, for different utility functions.
  • The direction $u$ in which an item is pushed can be selected in different ways. It may be chosen randomly (strategy 1) or as a normal to the item’s diameter, with a random choice of which side to push (strategy 2). However, the following method was primarily used: if the item is thin (with a diameter-to-width ratio greater than $3$), it is pushed to the left along the normal to its diameter; if the item is fat (ratio less than $3$), it is pushed to the right (strategy 3). Additionally, items may be pushed in the direction normal to their longest edge if they are fat (strategy 4). Figure 8 shows the average results over 10 runs for each instance from Table 1.
@@
Choosing the best parameters still remains a challenging task.
4 Engineering
We emphasize several critical points that must be carefully implemented in order to obtain an efficient code. In both geometric heuristics and local search, the majority of computational time is spent determining whether two items overlap.
4.1 Fast Crossing Detection
The items are preprocessed into data structures to minimize the computational time required to check whether two translated items intersect. Each item stores its axis-parallel bounding box. Given the items positions, if the translated bounding boxes do not overlap, further checks are unnecessary.
The data structure also includes a decomposition of the item boundary into monotone polylines, going from left to right (plus a boolean value telling whether the item is above or below the polyline). We call them chains. Each chain is also assigned its own bounding box. Convex chains have only two chains. Intersections between chains are computed in linear time. If the chains of two items cross, then the items overlap. In cases where the chains share common points without crossing (such as a shared vertex) special attention is required. These degenerate cases are particularly significant, as they correspond to configurations where the items are positioned to touch but not overlap, which is often the desired outcome.
4.2 The grid
When packing a large number of items, we employ a grid structure where each cell stores a list of items that intersect with it. To determine whether a new item can be placed at a specific position, we first identify the grid cells $C$ that the item crosses. From these cells, we retrieve the list $I_{C}$ of items intersecting them. To check if the new item $i$ can be placed at the considered position, we only need to test for intersections between $i$ and the items in $I_{C}$.
After optimizing the intersection tests, it is essential to also implement the push routine efficiently.
4.3 Push Routine and Issues
A key component of the greedy heuristic and optimization process is the operation that involves pushing an item $i$, initially placed at position $p$, in a chosen direction $u$ given by an integer vector. Although this algorithm may seem straightforward to implement, there are several pitfalls to avoid.
The push routine is designed to allow items to slide along the boundaries of other packed items and the container. To achieve this, we consider multiple integer vectors $v$ such that the dot product $u\cdot v$ is strictly positive, then translate the item as far as possible along the ray passing through $p$ in direction $v$. If a new valid position is found, we update the position $p$ and test another direction $v$. After testing a few directions $v$ without successfully moving the item $i$, we conclude the push routine.
We first choose the integer vector $u$, for example a normal vector to the diameter of the item. Let $u^{\prime}$ be a vector perpendicular to $u$. We try vectors $v$ of the form $v=u+\alpha u^{\prime}$ for integer $\alpha$ from $-8$ to $8$.
The primary computation involves translating the item in the direction of $v$. Here, we distinguish between pushing, which allows sliding in other directions as long as the dot product $u\cdot v$ increases, and translating, which refers to movement strictly in direction $v$.
The translation in the direction $v$ starts by computing the smallest factor $2^{k}$ for integer $k$, such that $p+2^{k}v$ lies outside the container. We then repeat the following procedure. We decrement $k$ until we find a position $p+2^{k}v$ where the item $i$ can be placed or $k$ becomes 00. If such a position is found, we update $p$ to the new position $p+2^{k}v$. We then repeat the procedure for the same vector $v$ and smaller values of $k$ until $p+v$ is not a valid position.
A problem that may arise is that, since $v$ is an integer vector, its length may be large, potentially leaving empty space between $p$ and $p+v$. To address this, we continue the process by investigating new positions along $p+2^{k}v$ for progressively smaller values of $k$ (e.g., $k=-1$, $k=-2$, and so on). Since only integer translations are allowed, we may need to round the coordinates of $2^{k}v$ when $k$ is negative. This introduces two potentially significant issues that need to be addressed. Both issue arise from the fact that division and rounding can alter the direction of the vector.
The first issue happens because $\textrm{round}(2^{k}v)$ is sometimes valid for much higher values of $k$, which takes a very long time. To solve this, we set $v$ to $\textrm{round}(2^{k}v)$. As the direction is slightly altered, we recompute the smallest integer $k$, such that $p+2^{k}v$ lies outside the container and start decreasing it as before.
The second issue is that $u\cdot\textrm{round}(2^{k}v)$ may become negative. If that happens, we stop the calculation and test other different directions $v$. Despite its apparent complexity, the push routine we implemented, which proceeds by discrete jumps and subsequently only calls the overlapping test, is surprisingly efficient.
5 Conclusion
The knapsack polygonal packing problem is extremely complex. In our experience competing in six annual CG:SHOP challenges, finding solutions that we believe are optimal has never been as hard. After months of work, we still managed to improve several of the smallest competition instances. Comparing the top three teams, there are only $4$ of $180$ instances for which at least two teams were in a tie with the best solution score.
The algorithms presented in this paper are the main ones that we used during the Challenge. With time to step back, these solutions hold numerous opportunities for improvement. There is significant room for enhancement, both in the clustering preprocessing step and in the initialization of a solution. Using clusters generation to select the items placed in each cell of the integer programming approach for large instances is another promising avenue for improvement. Additionally, our local search could be further refined by incorporating more sophisticated search techniques.
A high-level analysis of our results reveals that, for small instances, the combinatorial framework of the integer programming algorithm outperforms the greedy geometric approach. However, as the instances grow larger, the situation reverses. The combinatorial nature of the solutions demands increasingly high computational power, whereas the geometric approach offers fast heuristics that yield solutions of comparable quality. Better algorithms could be developed by combining the strengths of both approaches.